# Strong Coupling Constant Determination from the new CTEQ-TEA Global QCD Analysis

### v1

@@ IV.4.1 Reclustering algorithm
Focusing first on reclustering of $D^{2}$ into $N_{cl}$ clusters with equal numbers of data points (modulo $N_{cl}$), we randomly assign the $\vec{v}_{r}$ components into these clusters using a random generator engine (e.g., Mersenne-Twister [73]). We keep the same assignment of the residuals into clusters for all $\alpha_{s}(M_{Z})$ values on the grid. The $\chi^{2}_{E}$ dependence on $\alpha_{s}(M_{Z})$ for the $N_{cl}$ pseudoexperiments now defines the DT uncertainty on $\alpha_{s}(M_{Z})$. To construct the DT, we follow the same procedure as for the published data sets in Sec. IV.1. Namely, for each cluster $E$ we compute the quantiles at our CL, determine $\alpha_{s,min}^{(E)}$ and $\alpha_{s,max}^{(E)}$ for this quantile, and then construct the DT bounds as in Eq. (15), where $E$ now runs over the clusters.
For every $N_{cl}$, we repeat this procedure many (e.g., >200>200) times, each time reshuffling the assignment of $r_{0,k}^{(E)}$. Since the best-fit residuals and nuisance parameters enter $\chi^{2}_{R}$ as independent degrees-of-freedom, we can analogously shuffle the nuisance parameters from $R^{2}$ into random clusters; or, since the contributions from $R^{2}$ can only strengthen the DT uncertainties based solely on the residuals, we can initially omit $R^{2}$ for simplicity.
IV.4.2 Separate reclustering of residuals and nuisance parameters
We illustrate the possibility of separate reclustering of the residuals and nuisance parameters by two figures. Figure 9 goes back to the example of the GT and DT uncertainties for the published data sets in Fig. 6 for the CT25 baseline and now shows the DT uncertainties constructed only with $D^{2}$ and only with $R^{2}$ for the CT25prel baseline. [^5] These uncertainties do not change much between the two baselines. In this and following figure, the lower bands saturate at $\alpha_{s}(M_{Z})=0.113$, the lowest point on our $\alpha_{s}$ grid. The role of $R^{2}$ appears to be somewhat different for the GT and DT cases. In the GT case, the $R^{2}$ has a more constraining power on the uncertainty: if it were to be the only metric for the uncertainty, we would obtain a very small error (red dashed band). Conversely, without the $R^{2}$ contribution, the GT error (red solid lines in Fig. 9) are signicantly wider than the GT error based on both residuals and nuisance parameters, or $D^{2}+R^{2}$, in Fig. 6. However, $R^{2}$ is in general not perfectly determined, due to lack of full control on the correlated systematic uncertainties.
In the case of the DT, however, the errors without $R^{2}$ (blue dashed lines in Fig. 9) largely follow those based on $D^{2}+R^{2}$ in Fig. 6. In other words, the residuals appear to play a more important role in the DT case.
Figure 8 showed how the 68% and 90% CL error estimates change when we vary the DT clustering procesure. Figure 10 extends this comparison to show error estimates for other credibility levels, including the residuals only (that is, neglecting $R^{2}$). In this figure, the “Global T: no $R^{2}$” and “Dynamical T: no $R^{2}$” refer to the GT and DT uncertainties based on the full baseline. The “52 alt. clusters” and other lines refer to randomly generated clusters constructed by reshuffling the residuals in the 52 experimental data sets of CT25prel. For the alternative clusters, the lines show the envelopes of the DT estimates for 200 shuffles of the corresponding alternative clusters.
In an ideal scenario, in the absence of tensions among the data, alternative cluster configurations would produce approximately the same estimate for the $\alpha_{s}(M_{Z})$ uncertainty. In reality, when we partition into a lot of clusters, taking the outcome of the DT without $R^{2}$ by itself may already produce a stringent (likely, too stringent) uncertainty on $\alpha_{s}(M_{Z})$. However, the constraints may be weaker for other cluster configurations, and including $R^{2}$ into clustering may additionally reduce or increase the strength of the DT constraints by modifying the tensions, similarly to what we saw for the DT that reclustered the published data sets.
IV.4.3 $K$-folding impact on $\alpha_{s}$
In our next example, we will explore how the uncertainties on $\alpha_{s}(M_{Z})$ obtained in MC/ML approaches may be affected by common cross validation techniques that partition the baseline data set into training, testing, and occasionally validating components. An example of such procedure is $K$-folding [74, 75] introduced in the NNPDF determination [24] of $\alpha_{s}(M_{Z})$ to mitigate overfitting of the neural-network model for PDFs. Since $K$-folding involves reclustering of the baseline data and may be sensitive to data tensions, we will briefly review its possible impacts on the $\alpha_{s}(M_{Z})$ uncertainty.
cluster 0 cluster 1 cluster 2 cluster 3
$N_{\textrm{pt}}$ 1113 1113 1112 1112
    $\chi^{2}_{R}$ (no $R^{2}$)
shuffle 1 1094 1157 1281 1177
shuffle 2 1102 1158 1252 1198
shuffle 3 1065 1239 1251 1156
Table 2: Numbers of points and $\chi^{2}_{R}$ (without the systematic contribution $R^{2}$) for three random combinations (shuffles) of the residuals into four pseudodata clusters.
@@ IV.4.3 $K$-folding impact on $\alpha_{s}$
In the common $K$-folding procedure, each replica of the data baseline is partitioned into several folds (clusters), with one fold set aside as the test data set. The model is then trained on the union of remaining folds. A figure of merit called “loss function” $L$ is optimized so as to achieve good agreement with both the training and testing sets. The training procedure is iterated by excluding every fold in turn and training on the rest of the folds. In the $\alpha_{s}$ study, each iteration of such training+cross validation also returns the best-fit $\alpha_{s}(M_{Z})$ and its uncertainty. The final $\alpha_{s}(M_{Z})$ is determined by weighted averaging over all replicas and $K$-folding iterations. The weights may be given in terms of the respective $\alpha_{s}$ uncertainties from the iterations.
@@
Based on our preceding discussion, we expect the outcome of $K$-folding for $\alpha_{s}$ determination from the global data set to vary depending on the clustering algorithm and uncertainty definition. In the NNPDF study, each replica of the published data sets is partitioned into the training set and testing fold as 3:1, without breaking the data sets apart. The $\alpha_{s}$ uncertainty from each iteration of training is determined essentially according to the $\Delta\chi^{2}=1$ criterion. The final $\alpha_{s}(M_{Z})$ is obtained by averaging over the data replicas and $K$-folding iterations using $1/(\Delta\alpha_{s}(M_{Z}))^{2}$ from each iteration as the weight. Their resulting value $\alpha_{s}(M_{Z})=0.1194^{+0.0007}_{-0.0014}$. We pointed out in Sec. III.2 that the NNPDF central value changes depending on the normalization for the experimental systematic errors, in accord with our own findings. The NNPDF uncertainties are generally smaller than ours. While reproducing them is beyond the scope of our study, we can elucidate the variability of the preferred central values and their uncertainty under different clustering methods using the following simplified exercise.
@@
In contrast to the full $K$-folding procedure, we will not refit (optimize) the PDFs for each data partition, which would be too computationally expensive. Instead, we will examine uncertainties by reclustering the output of the CT25prel global fit, similarly to what was done in the previous sections. In addition, to assess their independent impact, we will separately consider the residual and nuisance-parameter contributions.
Dependence of the central $\alpha_{s}(M_{Z})$ on clustering. We generate multiple shuffles of the residuals into four clusters, with about 1113 points in each cluster, as outlined in Sec. IV.4.1. Table 2 reports the numbers of points and $\chi^{2}$ contributions for three representative shuffles in our series. Similarly, we can construct four $R^{2}$” clusters consisting of the shuffled best-fit nuisance parameters. Figure 11 illustrates the outcome of one such exercise, in which 100 instances of four clusters were generated. The figure separately depicts the individual $\chi^{2}/N_{pt}$ ratios for the residuals (purple lines) and nuisance-parameter(red-dashed lines). Each purple (red) line corresponds to a different shuffle configuration. The black lines represent the $\chi^{2}/N_{pt}$ average values. The stars represent the minima of the average curves which indicate the $\alpha_{s}(M_{Z})$ central-value preference. No tolerance criterion is applied here; we simply calculate the $\chi^{2}/N_{pt}$ in each cluster for many different random alternative configurations and varied $\alpha_{s}(M_{Z})$.
Shuffle 1
in 1-clusters: {0} {1} {3} {2}
& 3-clusters: {0,1,2}, {0,1,3}, {0,2,3} {1,2,3} N/A N/A
Shuffle 2
in 1-clusters: {0} {1} {3} {2}
& 3-clusters: {0,1,2}, {0,1,3}, {0,2,3} {1,2,3} N/A N/A
Shuffle 3
in 1-clusters: {0} {3} {1} {2}
& 3-clusters: {0,1,2}, {0,1,3}, {0,2,3} {1,2,3} N/A N/A
Table 3: For each shuffle of the $K$-folding exercise, we report the $\delta\alpha_{s}$ values found by applying the DT, as well as the one 1-cluster and three 3-clusters for which these $\delta\alpha_{s}$ are achieved. The columns containing the $\delta\alpha_{s}$ values and clusters are arranged so that the leftmost (rightmost) column contains the lowest (highest) $\delta\alpha_{s}$.
Just as one possibility, if the DT criterion is applied either to the “training set” consisting of 3 clusters or 1 “testing” cluster, it can render the uncertainty that can be substantially larger or smaller than the $^{+0.0007}_{-0.0014}$ from the NNPDF study. In its most granular realization, the DT criterion treats each residual as an independent “experiment”. Table 3 summarizes the outcome of applying such DT criterion to the three training clusters of the residuals (no $R^{2}$) and to the one left out. Here we show the DT $\alpha_{s}$ uncertainties for the same three shuffles as in Table 2. These uncertainties can be either substantially smaller or larger than the NNPDF one, depending on the specific partition. The table also highlights the observation that, within a combination of any three clusters, the DT uncertainty of the whole cluster is equal to the smallest individual DT uncertainty among the constituent clusters. For this specific example, cluster 0 has the lowest $\chi^{2}$ for the three shuffle configurations, cf. Table 2. When combined with the other two, this cluster drives the net uncertainty of the combination. For the fourth 3-cluster combination, the uncertainty is set by the second most precise individual cluster. According to the findings in Table 3, the uncertainty on $\alpha_{s}(M_{Z})$ is not uniformly reproduced by each test cluster, and this is mainly ascribed to tensions in the data. The results do not change if we consider the average of the errors over a large number of shuffle configurations.
This exercise can in principle be applied to any number of clusters with $N_{cl}\leq N_{pt}$ to explore the variability of $\alpha_{s}(M_{Z})$ and its uncertainty across the full granularity of data.
@@ V Bayesian Models for Uncertainty Estimation
The fundamental assumption of the tolerance prescriptions is that all experiments agree on the parameter of interest in a certain region, and this region can be estimated from the augmented loglikelihood by following the described procedures. In contrast, Bayesian model averaging assumes that some experiments are more likely to be correct than the others and estimates the underlying probabilities for them to be correct from the data. In this section, we describe how two such models can be used to determine the best-fit $\alpha_{s}$ and its uncertainty, working with the CT25 baseline data set as an example. As in the previous sections, we make a simplifying assumption that the correlations among the separate experimental data sets via the PDFs or other parameters are negligible, allowing us to treat these data sets as independent $\alpha_{s}$ determinations in the good approximation.
V.1 Bayesian Hierarchical Model
A Bayesian Hierarchical model (BHM) offers a method for combining $\alpha_{s}(M_{Z})$ determinations from the individual $\chi^{2}_{E}$ of each experiment by following the procedure outlined in Ref. [76]. According to this model, one introduces a distribution of hyperpriors, from which one can draw possible values of the parameter of interest, in our case $\alpha_{s}(M_{Z})$. Specifically, considering each $\alpha_{s_{i}}(M_{Z})$ to be the result of a random draw from a parent distribution, one can write
@@ V.1 Bayesian Hierarchical Model
where $p(\mu,\tau)$ is a hyperprior distribution for the hyperparameters $\mu$ and $\tau$. In Ref. [76] a $\mu$-independent hyperprior is proposed as follows:
$p(\tau)d\tau^{2}\propto\prod_{i=1}^{N}\left[\frac{1}{\sigma^{2}_{i}+\tau^{2}}\right]^{\frac{\beta}{2N}}d\tau^{2}\,$, (20)
where $\beta$ is a parameter that determines the nature of the prior being used, and $p(\tau)$ interpolates smoothly between a flat and a sharply peaked $\tau$ distribution. If $\beta=0$, one assumes a flat prior and hence a larger spread of the hyperprior distribution. Therefore, small values of $\beta$ are used for data sets that exhibit substantial disagreement, and vice-versa. We combine the data given in Table 4, according to which $N=3$ in Eq. (20) represents the number of $\alpha_{s}(M_{Z})$ measurements being combined and $\sigma_{i}$ are the corresponding uncertainties. For $\beta=0$, we find $\alpha_{s}(M_{Z})=0.1181^{+0.0030}_{-0.0031}$. The variation of the uncertainty estimate as a function of $\beta$ is shown in Fig. 13 (left), where we note that the $\alpha_{s}(M_{Z})$ uncertainty reduces as $\beta$ increases. Specifically, for $\beta=15$ we find $\alpha_{s}(M_{Z})=0.1182^{+0.0006}_{-0.0007}$. Given the large spread of $\alpha_{s}(M_{Z})$ values in Table 4 and as suggested in Ref. [76], we use smaller values of $\beta$. Rather than choosing a single value of $\beta$ we average over $0\leq\beta\leq 1$ which yields
@@
In the next section, we repeat this exercise using a variant of the implementation of the hierarchical model, where the choice of the hyperparameters is determined through the use of statistical information criteria.
Figure 13: Variation of the uncertainty on $\alpha_{s}(M_{Z})$ when using the BHM (left) with a hyperprior defined in Eq. (20), and when using the GMM (right) as a function of the error scaling factor $e_{S}$. The orange shaded region represents the uncertainty when setting $K=2$ in the GMM and the green shaded region represents the uncertainty for $K=1$, which corresponds to the usual $\chi^{2}$ fit. Also shown are vertical lines indicating the values of $e_{S}\approx 4.7$ and $e_{S}\approx 6.2$, respectively corresponding to the values of the PDG scale factor and the value of the global tolerance $T$, that would give the same uncertainty estimate as the GMM. For both plots, the data used in the combination is shown in Table 4.
V.2 Hierarchical Models via a Gaussian Mixture Model
In Ref. [77] a new Bayesian hierarchical model is proposed as a generalized likelihood, that is modeled via the Gaussian mixture model (GMM). In this section, we apply the GMM to the combination of the three $\alpha_{s}(M_{Z})$ measurements in Table 4, with the aim of capturing the multi-modal feature of this data set with large spread. [^6] A more complete implementation of the GMM requires modification of the $\chi^{2}$ loss function, contributed from each experimental data set included in the fit, that we do not perform here.
In the GMM, $K$ Gaussian components are introduced so that the likelihood is modeled as

### v2

@@ IV.4 Properties of the dynamic tolerance
$\textrm{Many clusters}:$ $\quad\textrm{DT}\to(\Delta\chi^{2}=1).$
When all experiments are combined, the dynamical tolerance reduces to the global one. When data are split into a large number of clusters (1000 in Fig. 8), the DT uncertainty essentially reduces to the $\Delta\chi^{2}=1$ uncertainty, when a single data point (degree of freedom) controls the allowed parameter displacement.
Figure 8: Global and dynamical tolerance uncertainties on $\alpha_{s}(M_{Z})$ at the 68% and 90% CL for various clustering procedures applied to the CT25prel baseline. DT uncertainties are estimated according to the default prescription, without and with including the sum $R^{2}$ of best-fit nuisance parameters, and optionally by reclustering the data into 3, 52, 100, and 1000 clusters as explained in Sec. IV.6. For the reclustered data, the DT errors are averaged over 200 alternative configurations.
IV.5 Data clustering safety of uncertainty estimates
The previous subsection points out that the DT uncertainty depends on how the data are clustered. The fitting groups typically construct the DT using the full published data sets as independent clusters, but other reasonable clustering choices are obviously possible.
Such sensitivity would raise the concern that the DT uncertainty estimate on a fundamental parameter is unsafe at some level with respect to the user-chosen data clustering prescription. In contrast, the GT is trivially clustering-safe.
The key point here is that reclustering of the data sets modifies the degree of tensions present in the data. For instance, in a version of the CT25prel fit to 52 data sets shown in Fig. 8, the GT estimate is $\alpha_{s}(M_{Z})=0.1179\pm 0.0024$, i.e. the error is practically symmetric. The default DT error, formed by the nominal clustering to the 52 data sets, is $0.1179^{+0.0024}_{-0.0011}$, showing a pronounced asymmetry. Upon reorganizing the data sets into 3 clusters, without breaking the data sets, in some cases the data sets are grouped so that their opposite pulls on $\alpha_{s}(M_{Z})$, and in other cases these pulls are maximally exacerbated. When the tensions are largely cancelled, the 3-cluster DT error becomes nearly symmetric and larger than the GT one, e.g., $0.1179^{+0.0032}_{-0.0028}$. When the clustering magnifies the tensions, the 3-cluster DT errors vary more and can be highly asymmetric or not, for instance, $0.1179^{+0.0021}_{-0.0008}$ or $0.1179^{+0.0016}_{-0.0019}$.
The DT estimates thus in principle should be supplemented by additional tests of the degree of their clustering (un)safety. In light of what was said above, the DT errors for the $\alpha_{s}$ determination vary between zero and the GT errors. One could take an envelope of these DT errors as a conservative estimate, or an average DT error over many clustering configurations as an optimal estimate. In our specific case, the latter estimate is close to the default one using the 52 individual experiments, and hence the final uncertainty combination will concentrate on the default DT tolerance.
IV.6 Data reclustering and K-folding
It is interesting to ask whether the tolerance-based methods can quantify the impact of tensions within the available data sets, not only among the data sets. In this subsection, we further investigate this question, which is generally complex but can be elucidated by partitioning the best-fit $\chi^{2}$ in Eq. (4) into alternative clusters of data residuals $r_{0,k}^{(E)}$ and nuisance parameters (NPs) $\lambda^{(E)}_{0,\alpha}$. This exercise reveals scaling properties of uncertainty estimates with respect to the number of clusters, although the scaling would be further affected by correlations among $r_{0,k}^{(E)}$ that are arguably less impactful and have been neglected. In particular, the estimates for GT and DT based on the residuals only (without $R^{2}$, the contribution from NPs) are overestimated by assuming $N_{\textrm{pt}}$ degrees of freedom for the contribution $D^{2}$ of the residuals instead of the actual $N_{\textrm{pt}}-N_{\textrm{par}}$. Working in this framework, we will summarize the procedure employed to partition the published data sets into many clusters illustrated in Fig. 8. Then, we will comment on the potential impact of cross validation procedures in the Monte-Carlo-based approaches on the $\alpha_{s}$ uncertainty estimate. The reader interested in the Bayesian methods and final combination can proceed directly to Secs. V and VI.
IV.6.1 Reclustering algorithm
If we neglect correlations via the PDF parameters, as was done to justify the DT to start with, then $r_{0,k}^{(E)}]^{2}$ and $\lambda^{(E)}_{0,\alpha}]^{2}$ from all data sets can be separately and independently collected (“shuffled”) into alternative clusters of data points, and then the DT can be formed from such clusters.
Working with the CT25prel baseline, we start by organizing all $\chi_{R}^{2}$ components in Eq. (4) into two vectors $\mathbf{v}_{r}$ and $\mathbf{v}_{\lambda}$ as
$\mathbf{v}_{r}$ $=$ $\left\{r^{(1)}_{0,1},\dots r^{(1)}_{0,N_{\textrm{pt}}^{(1)}},\dots r^{(N_{E})}_{0,1},\dots,r^{(N_{E})}_{0,{N_{\textrm{pt}}^{(N_{E}})}}\right\}\,$, (20)
$\mathbf{v}_{\lambda}$ $=$ $\left\{\lambda^{(1)}_{0,1},\dots,\lambda^{(1)}_{0,N_{\lambda}^{(1)}},\dots,\lambda^{(N_{E})}_{0,1},\dots,\lambda^{(N_{E})}_{0,N_{\lambda}^{(N_{E})}}\right\}\,$, (21)
[…]
@@ IV.6.1 Reclustering algorithm
We randomly assign the $\mathbf{v}_{r}$ and $\mathbf{v}_{\lambda}$ components to $N_{cl}$ clusters with equal numbers of data points (modulo $N_{cl}$), keeping this assignment same for all $\alpha_{s}$. We construct the DT according to the same procedure as for the published data sets in Sec. IV.3. Namely, for each cluster $E$ we compute the $\chi^{2}_{E}$ quantile at our CL, determine $\alpha_{s,min}^{(E)}$ and $\alpha_{s,max}^{(E)}$ for this quantile, and then construct the DT bounds as in Eq. (15), where $E$ now runs over the clusters. We examine the $\alpha_{s}$ dependence of DT by (a) constructing the DT interval for each of $N_{\alpha_{s}}=21$ values of $\alpha_{s}(M_{Z})$ on a grid between 0.113 and 0.129, and (b) repeating this procedure many (e.g., >200>200) times, each time reshuffling the assignment of $r_{0,k}^{(E)}$ and/or $\lambda^{(E)}_{0,\alpha}$. We can compute DT for $D^{2}$ only, $R^{2}$ only, or the whole $D^{2}+R^{2}$. We can do this because, in our approximation, the best-fit residuals and nuisance parameters enter $\chi^{2}_{R}$ as independent degrees of freedom.
In Fig. 7, the first five lines compare the GT and DT uncertainties at 68% and 90% CL computed based on the full $\chi^{2}_{E}$, residuals ($D^{2}$) only, and nuisance parameters ($R^{2}$) only. This comparison reveals elevated tension in the $R^{2}$, causing the DT uncertainty based on $R^{2}$ only (line 5) to be smaller than those based on the full $D^{2}+R^{2}$ (line 2) and $D^{2}$ only (line 4). Appendix B further explores the GT and DT estimates collected in Fig. 7, also providing them at other credibility levels.
IV.6.2 K-folding impact on $\alpha_{s}$
Cross validation methods such as jackknife resampling [82, 83, 84] are widely adopted to estimate the data selection uncertainty. In a classical realization, the parameter of interest is first estimated from a part (fold) of the baseline data; the parameter’s expectation value and variance are determined by averaging over multiple random selections of such folds. Cross validation is also a part of MC/ML approaches and involves partitioning the baseline data set into training, testing, and occasionally validating components. One can ask, for instance, to what extent the cross validation technique of “K-folding” deployed in the recent NNPDF determination [24] of $\alpha_{s}(M_{Z})$ may be sensitive to data tensions. Since K-folding involves reclustering of the baseline data, we will briefly review its possible impacts on the $\alpha_{s}(M_{Z})$ uncertainty.
@@ IV.6.2 K-folding impact on $\alpha_{s}$
In the common K-folding procedure, each replica of the data baseline is partitioned into several folds (clusters), with one fold set aside as the test data set. The model is then trained on the union of remaining folds. A figure of merit called a “loss function” $L$ is optimized so as to achieve good agreement with both the training and testing sets. The training procedure is iterated by excluding every fold in turn and training on the rest of the folds. In the $\alpha_{s}$ study, each such iteration also returns the best-fit $\alpha_{s}(M_{Z})$ and its uncertainty. The final $\alpha_{s}(M_{Z})$ is determined by weighted averaging over all replicas and K-folding iterations. The weights may be given in terms of the respective $\alpha_{s}$ uncertainties from the iterations.
@@
In the NNPDF study [24], each replica of the published data sets is partitioned into the training set and testing fold as 3:1, without breaking the data sets apart. The $\alpha_{s}$ uncertainty from each iteration of training is determined essentially according to the $\Delta\chi^{2}=1$ criterion. The final $\alpha_{s}(M_{Z})$ is obtained by averaging over the data replicas and K-folding iterations using $1/(\Delta\alpha_{s}(M_{Z}))^{2}$ from each iteration as the weight. Their resulting value $\alpha_{s}(M_{Z})=0.1194^{+0.0007}_{-0.0014}$. The lower of these uncertainties is roughly equal to the quadrature sum of our $\Delta\chi^{2}=1$ and PDF parametrization uncertainties. The NNPDF uncertainty is smaller than our $\delta\alpha_{s}\approx 0.002$, plausibly because the latter also includes the data selection uncertainty.
We posit, based on the preceding discussion, that the reclustering involved in K-folding should affect $\delta\alpha_{s}$. For a reliable estimate, $\delta\alpha_{s}$ should be averaged over a representative ensemble of K-folding/cross-validation partitions.
While such an exhaustive analysis, especially refitting the PDFs for multiple data partitions, is beyond the scope of our article, Appendix C highlights several considerations based on the following simplified exercise.
@@
Working again with vector $\mathbf{v}_{r}$ of about 4000 best-fit residuals from the CT25prel baseline, we partition the residuals into four clusters of about 1000 members each. Similarly, we partition nuisance parameters from vector $\mathbf{v}_{\lambda}$ into four approximately equal clusters. We repeat this step 100 times, each time randomly modifying the partitions. For each partition (alternative configuration) and for the averages over 100 alternative configurations, we plot the $\alpha_{s}$ dependence of $\chi^{2}_{R}/N_{\textrm{pt}}$ in each cluster in Fig. 9. We make an important observation, namely, that the $\chi^{2}_{R}/N_{\textrm{pt}}$ distributions in Fig. 9 are flatter than would be expected from an ideal distribution of a cluster of $\approx$ 1000 made-up residuals that are consistently distributed according to $\mathcal{N}(0,1)$ around a single reference value of $\alpha_{s}(M_{Z})=0.119.$ Figure 10 shows this idealized distribution rendering a deeper parabola.
The flatter $\chi^{2}_{R}$ of the clusters of actual residuals in Fig. 9 indicates a larger uncertainty than $\approx 0.0003$ obtained by fitting the ideal parabola in Fig. 10, with the latter being in a perfect agreement with $\Delta\chi^{2}=1$. The second part of this exercise tries to estimate the actual uncertainty, keeping in mind that it depends on how the clusters are formed. To gain an insight about the practical K-folding studies, we determine $\delta\alpha_{s}$ from random combinations of three out of four above clusters. In this specific example, and using the DT, we find that $\delta\alpha_{s}$ varies between $\approx$ 0.0002 and 0.002 depending on the magnitude of tension in the dominant cluster of each configuration.
This exercise can be expanded in a more realistic cross-validation setup to explore the variability of $\alpha_{s}(M_{Z})$ and its uncertainty across the full granularity of data. Appendix C contains further details, including concrete uncertainty estimates.
@@ V Bayesian Models for Uncertainty Estimation
[…]
V.1 A Bayesian Hierarchical Model
In combining the different values of $\alpha_{s,i}(M_{Z})$ from multiple experiments, a key difficulty is that the quoted uncertainties from each experiment may not fully capture all sources of systematic error, theory assumptions, or hidden correlations. A naive combination, such as the simple weighted average in Sec. IV.2, implicitly assumes that all measurements are mutually consistent and that their reported uncertainties are complete. When this assumption fails, the resulting average can become artificially precise or biased toward experiments that happen to quote smaller, but possibly underestimated, error bars.
A Bayesian hierarchical model (BHM) provides a principled way to address this problem. Instead of treating each experimental determination $\alpha_{s_{i}}(M_{Z})$ as an estimate of a single common parameter with known variance, the hierarchical approach acknowledges the possibility of unmodeled spread among the measurements by introducing hyperparameters that describe a latent “parent” distribution from which the individual measurements of $\alpha_{s,i}(M_{Z})$ are drawn. Specifically, considering each $\alpha_{s_{i}}(M_{Z})$ to be the result of a random draw from a parent distribution, one can write
@@ V.1 A Bayesian Hierarchical Model
where $p(\mu,\tau)$ is a hyperprior distribution for the hyperparameters $\mu$ and $\tau$.
Before introducing the specific hierarchical model used in this analysis, it is important to comment on the choice of hyperprior distribution. In Bayesian hierarchical methods, the role of the hyperprior is not to encode detailed knowledge about the data, but rather to parameterize our uncertainty about the possible degree of inter-experimental variation. While a fully data-driven determination of the hyperprior is in principle possible, it generally requires a more complex model structure. We consider two possible models, one where the the choice of hyperprior is well-motivated but not derived from data (in this subsection), and the other where the hyperprior is determined through data-driven methods (in Sec. V.2).
A suitable guiding principle in choosing the hyperprior is that it should (i) be weakly informative, so that it does not dominate the posterior when the data are mutually consistent, and (ii) be flexible enough to accommodate additional dispersion among the measurements when they exhibit tension. The form proposed in Ref. [85] satisfies these requirements while avoiding ad hoc error inflation or outlier removal. It introduces a single hyperparameter $\tau$ describing the possible extra variance beyond the quoted experimental uncertainties, and assigns to it a $\mu$-independent hyperprior,
$p(\tau)\,d\tau^{2}\propto\prod_{i=1}^{N}\left[\frac{1}{\sigma_{i}^{2}+\tau^{2}}\right]^{\beta/(2N)}d\tau^{2},$ (24)
which smoothly interpolates between a flat prior ($\beta=0$) and a sharply peaked one for larger values of $\beta$. This construction allows the analyst to encode a controlled level of prior belief about the reliability of the quoted uncertainties: smaller values of $\beta$ correspond to a more permissive prior, suitable when the data exhibit significant disagreement, while larger values correspond to the assumption that most of the quoted uncertainties are reliable. This model was introduced as a more general alternative to the PDG prescription [85].
We combine the data given in Table 2, according to which $N=3$ in Eq. (24) represents the number of $\alpha_{s}(M_{Z})$ measurements being combined, and $\sigma_{i}$ are the corresponding uncertainties. For $\beta=0$, we find $\alpha_{s}(M_{Z})=0.1181^{+0.0030}_{-0.0031}$. The variation of the uncertainty estimate as a function of $\beta$ is shown in Fig. 11 (left), where we note that the $\alpha_{s}(M_{Z})$ uncertainty reduces as $\beta$ increases. Specifically, for $\beta=15$ we find $\alpha_{s}(M_{Z})=0.1182^{+0.0006}_{-0.0007}$. Given the large spread of $\alpha_{s}(M_{Z})$ values in Table 2, and following the suggestion of Ref. [85], we use smaller values of $\beta$. Rather than choosing a single value of $\beta$, we average over $0\leq\beta\leq 1$, which yields
@@
Figure 11: Variation of the uncertainty on $\alpha_{s}(M_{Z})$ when using the BHM (left) with a hyperprior defined in Eq. (24), and when using the GMM (right) as a function of the error scaling factor $e_{S}$. The orange shaded region represents the uncertainty when setting $K=2$ in the GMM, and the green shaded region represents the uncertainty for $K=1$, which corresponds to the usual $\chi^{2}$ fit. Also shown are vertical lines indicating the values of $e_{S}\approx 4.7$ and $e_{S}\approx 6.2$, respectively corresponding to the PDG scale factor in Eq. (13) and tolerance $T$ with standard combination ($K=1$) that would give the same uncertainty estimate as the GMM with $K=2$. For both plots, the data used in the combination is shown in Table 2.
V.2 A Gaussian Mixture Model
The Gaussian mixture model (GMM) proposed in Ref. [86] is a new Bayesian hierarchical model in which the hyperprior is determined from data through the use of statistical information criteria. In this section, we summarize an application of the GMM to combine the $\alpha_{s}(M_{Z})$ values with a large spread in Table 2 by capturing the multimodal feature of these data. Appendix D discusses technical aspects of this combination.
In the GMM, $K$ Gaussian components are introduced, so that the likelihood is modeled as